OpenStudy (anonymous):

how do i find the zero of f(x)=5x-2?

OpenStudy (whpalmer4):

Finding the zero(s) of f(x) is simply finding the values of x where f(x)=0. \[f(x) = 5x -2 = 0\]Solve for x.

OpenStudy (anonymous):

5x - 2 =0 +5 and +2 7=0 i think thats right but im not for sure

OpenStudy (anonymous):

sorry i diddnt know for sure

OpenStudy (anonymous):

is it multiple choice

OpenStudy (anonymous):